# Estimating the Number of Street Vendors in New York City††thanks: Thanks to Mohamed Attia, Eric Auerbach, Debipriya

### v1

@@ Abstract
We estimate the number of street vendors in New York City. We first summarize the process by which vendors receive licenses and permits to legally operate in New York City. We then describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Finally, we calculate the total number of vendors using ratio estimation. We find that approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also provide a theoretical justification of our estimates based on the theory of point processes and a discussion of their accuracy and implications. In particular, our estimates suggest the American Community Survey fails to cover the majority of New York City street vendors.
@@ 1. Introduction
In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in three sections. In Section 2, we review the process by which vendors receive permits and licenses, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit. In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We also provide a theoretical justification of our estimates based on the theory of point processes and a discussion of their accuracy. In Section 4, we conclude by briefly discussing some implications of our estimates. In particular, our estimates suggest the American Community Survey fails to cover the majority of New York City street vendors.
@@ 2. We consider two types of vendors: food vendors and merchandise vendors
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm.
@@ 2.1 The number of licenses and permits are limited by local law
A long list of rules determine the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of merchandise licenses and food permits are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.222Note that our estimate reflects the number of street vendors as of 2021, prior to the implementation of Local Law 18 of 2021, which was enacted concurrently. Local Law 18 made several changes to the license and permit process. These changes do not impact our estimates and therefore are not discussed in this paper.
There are several types of mobile food vendor permits with varying restrictions. Of the 5,100 permits available: 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
There are three relevant exceptions to these rules. The first exception is that merchandise vendors who are veterans are not subject to the limit of 853 licenses. According to data obtained by Mosher and Turnquist (2024), there are approximately 1,000 merchandise vendors who are veterans in New York City. We assume the number of veterans selling merchandise without a license is negligible.
@@ 2.2 The Street Vendor Project surveyed two thousand street vendors in 2021
The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors. SVP administered a survey to approximately 2,000 street vendors while distributing Coronavirus relief aid in 2021. The aid was a one time payment of $1,000, available to any individual that owned, operated, or was otherwise employed by a street vending business in New York City between 2020 and 2021. First Amendment vendors were also eligible. There were no limits based on the size of the business, the number of sales, or whether licensed and permitted. All individuals eligible for aid were invited to complete a survey.
@@
The population of street vendors estimated in this paper is the population eligible for aid (as determined by SVP) that self-identifies as either a food or merchandise vendor. SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods. Survey operations continued until $2,415,000 in funds were distributed, yielding 2,060 responses. With limited exceptions, every vendor that received aid participated in the survey.
@@
The survey included 100 questions and was conducted in eight different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof.333Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English. Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl or Quechua, often spoke Spanish with a high level of fluency. The survey items solicited a variety of information from vendors, such as logistical information—e.g., vending location, residential location, and frequency of operation—economic information—e.g., items sold, income, and expenses—and demographic information—e.g., age, race, ethnicity, and immigration status. Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
@@
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents are listed by neighborhood in Section 6.1, Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 6.2, Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are excluded from Table 1 and Figure 1, although doing so does not meaningfully change the estimated number of vendors that operate in New York City. Veteran merchandise vendors are also excluded from Table 1 and Figure 1, although these vendors are reflected in our city-wide estimate of 23,000 vendors.
@@ 3. We calculate the number of street vendors using ratio estimation
In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership. In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data. In Section 3.3, we discuss the accuracy and implications of our estimates. Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 6.3.
@@ 3.1 We estimate 23,000 street vendors operate in New York City
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for a modern reference. The popularity of ratio estimation is due in large part to its simplicity: Among the food vendors surveyed, roughly a quarter had a permit. Suppose this ratio is representative in that one quarter of all food vendors have a permit. It follows that the total number of food vendors would be because one quarter of is approximately .
@@
The main assumption of ratio estimation is that the survey data are representative in the sense that, were survey operations to continue, the percentage of respondents with a permit (in the case of food vendors) or a license (in the case of merchandise vendors) would remain the same. If we further assume that the percentage does not vary systematically by location, such that the spatial distribution of respondents is representative, we can further estimate the total number of vendors by neighborhood.444Ratio estimation by neighborhood is not possible since only the locations of the vendors who participated in the survey are known. These estimates are listed in Section 6.1, Table 1. We find that approximately one quarter of street vendors are located in the Chelsea and Clinton neighborhoods of Manhattan. Another quarter is located in West Queens and Lower Manhattan.
We check our estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on Friday, May 13, 2022, between 4 - 8 p.m. and Sunday, May 15, 2022, between 12 - 7 p.m., in which the number of visible street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors in the Fordham area (Fordham Road BID and Street Vendor Project 2024), whereas Table 1 indicates 206 vendors in Bronx Park and Fordham—an estimate approximately ten percent higher. To demonstrate that this difference can be explained by sampling variation, we derive the standard error of the ratio estimate in Section 3.2.
@@ 3.2 We model the location of street vendors as a Poisson process
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to examine the underlying assumptions and characterize the accuracy of the estimate.
Let denote a family of spatial Poisson processes, referencing the location of each street vending establishment (cart, stall, truck, etc.) at the time the survey was conducted. The index denotes the status of the establishment. To fix ideas, let denote mobile food vending establishments without a permit, and let denote mobile food vending establishments with a permit.
The Poisson process assumption may be justified by the law of rare events. New York City can be partitioned into a large number of theoretically vendable locations. Whether an establisment is located within a partition has a vanishingly small probability, such that the number of establishments in any area is well approximated by a Poisson distribution. Note that the approximation is accurate even if these probabilities are weakly dependent. e.g., See Freedman (1974), Chen (1975), and Serfling (1975).
We are primarily interested in the number of vendors, not the number of vending establishments. We assume that at the time of the survey, each establishment is associated with a random number of individuals, distributed independently such that by the Marking Theorem (Kingman 1992), the locations of individuals are represented by a family of spatial Poisson processes with mean measures . It follows that the number of individuals, , is distributed Poisson with mean , i.e.,
We further assume an individual responds to the survey with probability such that by Campbell’s Theorem (Kingman 1992), the number of respondents is distributed Poisson with mean . i.e.,
@@
It is easy to verify that the maximum likelihood estimates for and are
@@ 3.3 The Street Vendor Project survey is spatially representative
We compare the spatial distribution of the Street Vendor Project survey respondents to the spatial distribution of street vendors suggested by administrative and federal survey data. The primary purpose is to assess the main assumption described in Section 3.1 and 3.2: that the response probability is spatially uncorrelated with the mean measure.
We think this assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid. We argue that trust in the Project and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects personal circumstances that are largely unrelated to location.
@@
These comparisons are visualized in Section 6.2, Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels), although there are several small discrepancies. For example, the OATH data suggest a larger percentage of vendors work in South Brooklyn (Coney Island and Brighton Beach) than captured by the Street Vendor Project survey, and the ACS data suggest a larger percentage of street vendors live in South Queens (Rockaway). These differences may reflect the exceptions noted in Section 3.2. Even so, both the OATH and ACS data suggest these locations reflect only a small percentage of New York City vendors, and thus an undercount is likely to have little impact on our overall estimate.
@@
We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated. That is because other areas may be overcounted (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate. Note that in comparing the residencies of vendors, we assume that the population of vendors who work outside of New York City but reside within New York City is negligible.

### v2

@@ Abstract
We estimate the number of street vendors in New York City. First, we summarize the process by which vendors receive licenses and permits to operate legally in New York City. Second, we describe a survey that was administered by the Street Vendor Project while distributing Coronavirus relief aid to vendors operating in New York City both with and without a license or permit. Third, we calculate the total number of vendors using ratio estimation, finding approximately 23,000 street vendors operate in New York City—20,500 mobile food vendors and 2,300 general merchandise vendors—with one third located in just six ZIP Codes—11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. Fourth, we provide a theoretical justification of our estimates based on the theory of point processes. Finally, we evaluate the accuracy of the ratio estimator when the distribution of vendors is explained by a Poisson or Yule process, and we discuss several policy implications. In particular, our estimates suggest the American Community Survey misses the majority of New York City street vendors.
@@ 1. Introduction
In this paper, we propose an estimate for the number of street vendors in New York City, including those that operate without a license or permit. We present our work in four sections. In Section 2, we review the process by which vendors receive permits and licenses, and we describe a survey administered by the Street Vendor Project at the Urban Justice Center while distributing Coronavirus relief aid to vendors operating both with and without a license or permit.
In Section 3, we use the fact that the number of licenses and permits are limited by law to construct a ratio estimator for the number of street vendors. We find approximately 23,000 street vendors operate in New York City: 20,500 mobile food vendors and 2,300 general merchandise vendors. One third are located in just six ZIP Codes: 11368 (16%), 11372 (3%), and 11354 (3%) in North and West Queens and 10036 (5%), 10019 (4%), and 10001 (3%) in the Chelsea and Clinton neighborhoods of Manhattan. We then provide a theoretical justification of our estimates, assuming the spatial distribution of survey respondents is representative and well-approximated by a family of inhomogeneous Poisson processes.
In Section 4, we examine the validity of these assumptions. We evaluate whether the spatial distribution of respondents is representative by comparing it to the American Community Survey and administrative data, and we evaluate the Poisson process assumption by considering a more general model in which the arrival of street vendors follows a pure birth process, of which the Poisson and Yule processes are special cases.
In Section 5, we discuss some implications of our estimates. In particular, we find that though the spatial distribution of respondents matches the results of the American Community Survey, the two disagree on the size of New York City’s street vending population. Our estimates suggest the American Community Survey misses the majority of New York City street vendors.
@@ 2. We consider two types of vendors: food vendors and merchandise vendors
A street vendor is any individual who sells goods from a mobile vending unit instead of a store. We distinguish between the mobile vending unit from which the goods are sold (i.e., the vending establishment) and the individuals who own and/or operate that unit (i.e., the vendors). Note that an establishment refers to a single vending unit. There may be multiple vendors associated with any establishment, and multiple establishments may be owned by a single individual or firm.222For example, suppose Sally owns the business Sally’s Salads, which consists of two food trucks, each staffed by three employees. Then Sally’s Salads counts as one business, two establishments, and seven vendors.
@@ 2.1 The number of licenses and permits are limited by local law
A long list of rules determines the individuals, locations, and time periods during which vendors can legally sell their goods in New York City. Most relevant is the fact that a merchandise vendor requires a license to operate legally, and a food vendor requires both a license and a permit. The number of food permits and merchandise licenses are limited by law: 5,100 permits and 853 general merchandise licenses are available. Mobile food vending licenses are not limited. Note that a mobile food vending permit is issued to an individual or business to allow for the sale of food from a specific mobile food vending unit, such as a cart or truck. A mobile food vending license authorizes an individual to prepare or serve food from a permitted mobile food vending unit.333Note that our estimate reflects the number of street vendors as of 2021, prior to the implementation of Local Law 18 of 2021, which was enacted concurrently. Local Law 18 made several changes to the license and permit process. These changes do not impact our estimates and therefore are not discussed in this paper.
There are several types of mobile food vendor permits with varying restrictions. Of the 5,100 permits available, 200 are borough permits that limit vendors to one of the five boroughs; 100 are reserved for veterans or vendors with a disability; 1,000 are seasonal and valid only from April to October; 1,000 are green cart permits that limit vendors to selling fruit, vegetables, plain nuts, and water; and 2,800 are unrestricted. Multiple individuals with food vendor licenses can legally operate from one permitted food vending unit. Merchandise licenses are renewed annually, while food licenses and permits are renewed biennially.
There are three relevant exceptions to these rules. The first exception is that merchandise vendors who are veterans are not subject to the limit of 853 licenses. According to data obtained by Mosher and Turnquist (2024), there are approximately 1,000 licensed merchandise vendors who are veterans in New York City. We assume the number of veterans selling merchandise without a license is negligible.
@@ 2.2 The Street Vendor Project surveyed two thousand street vendors in 2021
The Street Vendor Project (SVP), part of the Urban Justice Center, is a non-profit organization that advocates on behalf of New York City street vendors. SVP administered a survey to approximately 2,000 street vendors while distributing Coronavirus relief aid in 2021. The aid was a one-time payment of $1,000, available to any individual that owned, operated, or was otherwise employed by a street vending business in New York City between 2020 and 2021. First Amendment vendors were also eligible. There were no limits based on the size of the business, the number of sales, or whether licensed and/or permitted. All individuals eligible for aid were invited to complete a survey.
@@
The population of street vendors estimated in this paper is the population eligible for aid (as determined by SVP) that self-identifies as either a food or merchandise vendor. SVP found aid-eligible individuals through membership lists, referrals, and canvassing operations in which SVP affiliates visited neighborhoods. Survey operations continued until $2,415,000 in funds were distributed, yielding 2,060 responses.
@@
The survey included 100 questions and was conducted in eight different languages: Arabic, Bangla, Cantonese, English, French, Mandarin, Spanish, Tibetan, and Wolof.444Vendors who spoke other languages, such as Dari, Farsi, Russian, and Turkish, typically spoke English with a high level of fluency, and those surveys were conducted in English. Similarly, vendors who spoke indigenous languages from Latin America, such as Nahuatl or Quechua, often spoke Spanish with a high level of fluency. The survey items solicited a variety of information from vendors, such as logistical information (e.g., vending location, residential location, and frequency of operation), economic information (e.g., items sold, income, and expenses), and demographic information (e.g., age, race, ethnicity, and immigration status). Most relevant is the fact that respondents classified themselves by the goods they sold (e.g., food vendors, merchandise vendors, First Amendment vendors, etc.), and respondents indicated whether they had the relevant licenses and permits to vend. Of the 2,060 responses, 1,400 identified as food vendors and 559 as merchandise vendors. The remaining 101 respondents were predominantly First Amendment Vendors, which we exclude from our analysis.
@@
Of the 1,400 food vendors, 349 (25%) indicated they had a permit to vend. Of the 559 merchandise vendors, 505 were not veterans, of which 308 (61%) indicated they had a license to vend. The number of respondents is listed by neighborhood in Section 7.1, Table 1. (Neighborhoods with few respondents are grouped together.) A map of the number of respondents by ZIP Code Tabulation Area (ZIP Code) is shown in the top left panel of Section 7.2, Figure 1. Approximately five percent of vendors did not identify their vending location. These vendors are excluded from Table 1 and Figure 1, although doing so does not meaningfully change the estimated number of vendors that operate in New York City. Veteran merchandise vendors are also excluded from Table 1 and Figure 1, although these vendors are reflected in our city-wide estimate of 23,000 vendors.
@@ 3. We calculate the number of street vendors using ratio estimation
In Section 3.1, we provide a simple explanation of ratio estimation, which is intended to be accessible to a general readership. In Section 3.2, we provide a theoretical justification of our estimates based on the theory of point processes, which, while more technical, is intended to provide a somewhat general argument for ratio estimation with spatial data when a design-based approach cannot be justified. Inference is conducted using increasing domain asymptotics, and the details are outlined in Section 7.3.
@@ 3.1 We estimate 23,000 street vendors operate in New York City
Ratio estimation is a common approach for estimating the size of a population. See Cochran (1978) and Hald (1998, chap. 16) for a historical discussion and Lohr (2021, chap. 4) for an introduction. The popularity of ratio estimation is due in large part to its simplicity: Among the food vendors surveyed, roughly a quarter had a permit. Suppose this ratio is representative in that one quarter of all food vendors have a permit. It follows that the total number of food vendors is because one quarter of is approximately , the number of permits available.
@@
The main assumption of ratio estimation is that the survey data are representative in the sense that, were survey operations to continue, the percentage of respondents with a permit (in the case of food vendors) or a license (in the case of merchandise vendors) would remain the same. If we further assume that the spatial distribution of respondents is representative, we can also estimate the total number of vendors by neighborhood.555We use the same ratio for each neighborhood. Ratio estimation by neighborhood is not possible since only the locations of the vendors who participated in the survey are known. These estimates are listed in Section 7.1, Table 1. We find that approximately one quarter of street vendors are located in the Chelsea and Clinton neighborhoods of Manhattan. Another quarter is located in West Queens and Lower Manhattan.
We check our estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of visible street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors in the Fordham area (Fordham Road BID and Street Vendor Project 2024), whereas Table 1 indicates 206 vendors in Bronx Park and Fordham—an estimate approximately ten percent higher. To demonstrate that this difference can be explained by sampling variation, we derive the standard error of the ratio estimate in Section 3.2.
@@ 3.2 We model the location of street vendors as a Poisson process
We provide a point process justification of the ratio estimation procedure described in Section 3.1. The purpose is to examine the underlying assumptions and derive the standard error of the estimate.
Let denote a family of inhomogeneous spatial Poisson processes, referencing the location of each street vendor at the time the survey was conducted. The index denotes the status of the vendor. To fix ideas, let indicate mobile food vendors with a permit, and let indicate mobile food vendors without a permit.
The Poisson process assumption may be justified by the law of rare events. New York City can be partitioned into a large number of theoretically vendable locations. Whether a vendor is located within a partition has a vanishingly small probability, such that the number of vendors in any area is well approximated by a Poisson distribution. This approximation is accurate even if these probabilities are weakly dependent. For example, see Freedman (1974), Chen (1975), and Serfling (1975). Note that other justifications are possible. For example, see Section 4.2 for a justification of the Poisson process based on the theory of birth processes.
Let denote the mean measures of so that the number of individuals in any region , , is distributed Poisson with mean , i.e.,
We assume an individual responds to the survey with probability such that by Campbell’s Theorem (Kingman 1992), the number of respondents is distributed Poisson with mean . i.e.,
@@
The maximum likelihood estimates for and are
@@
4. We examine the validity of our assumptions using additional information
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is consistent. A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. This assumption is necessary to derive the standard errors.
In this section, we examine each assumption. Since neither can be evaluated from the survey data alone, we consider additional information. In Section 4.1, we compare the spatial distribution of the respondents to data from the American Community Survey Public Use Micro Sample 2017-2021 and the New York City Office of Administrative Trials and Hearings. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.666The results of this section can also be derived from a spatial Negative Binomial Lévy Process in a manner that mirrors Section 3.2. The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
@@ 4.1 The Street Vendor Project survey is spatially representative
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative (i.e., the response probability is spatially uncorrelated with the mean measure). If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem.
If this assumption is violated, then may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio would not be representative, and would likely underestimate even in large samples.
We think this assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid. We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
@@
These comparisons are visualized in Section 7.2, Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels), although there are several small discrepancies. For example, the OATH data suggest a larger percentage of vendors work in South Brooklyn (Coney Island and Brighton Beach) than captured by the Street Vendor Project survey, and the ACS data suggest a larger percentage of street vendors live in South Queens (Rockaway). However, both the OATH and ACS data suggest these locations reflect only a small percentage of New York City vendors, and thus any resulting undercount is likely to have little impact on our overall estimate.
@@
We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated. That is because this undercount may be compensated by an overcount in other areas (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
4.2 We model the arrival of street vendors as a pure birth process
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because the consistency of holds under the wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated.777The ratio estimator may no longer be asymptotically efficient as well. i.e., In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies the ratio estimator is asymptotically efficient. However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
We think the Poisson assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2. Nevertheless, to evaluate how the standard errors might change when this assumption is violated, we consider a general model that describes how street vendors arrived at their present locations. Specifically, we consider two families of pure birth processes: one with a constant arrival rate (or “birth rate”) in which the Poisson assumption holds and one with a linear arrival rate in which the Poisson assumption does not hold.
Suppose vendors serve customers who congregate at fixed locations, which we call markets. Let denote the number of markets in area . If every markets is in a competitive equilibrium such that the expected profit for a new vendor is the same at each market, then the time between vendor arrives at each market may be well-approximated by an exponential distribution with a constant arrival rate .888Arriving vendors may form a new vending establishment or sell their labor to an existing establishment. Further suppose each market starts with one vendor at . Then at the time the survey was conducted (), the number of vendors, , would follow a Poisson distribution with rate , and the results of Section 3.2 hold.
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925). That is, let denote the number of vendors in area at time so that the time between vendor arrivals at each market follows an exponential distribution with rate . If we assume further that each market started with one vendor at , then the number of vendors at the time the survey was conducted (), , would follow a negative binomial distribution with mean and variance . i.e.,
where we parameterize the negative binomial distribution by its mean and variance. In this case, exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details.999Note that we define the negative binomial and negative hypergeometric distributions as the total number of trials required to get a predetermined number of successes. The negative binomial assumes sampling with replacement, while the negative hypergeometric distribution assumes sampling without replacement.
To derive the standard errors, suppose a portion of the markets, , were selected for the survey so that
The conditional distribution of given is then
with conditional mean and variance
so that under increasing domain asymptotics, the standard error of
is asymptotically normal with mean and standard error
where
See Section 7.3 for details.
Note that this standard error is identical to the standard error from Section 3.2 except for the weights and . These weights are the result of the extra-Poisson variation induced by preferential attachment—i.e., the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
The weights and correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that and are no more than 10 times larger than . It follows that and are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most be 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market only contains one or two vendors.